OpenStudy (mowrey):

How many perfect cubes are there between 100 and 10,000,000?

OpenStudy (mowrey):

100 * 100 * 100 = 100^3 = 10,000,000 4^3 = 64 5^3 = 125 99^3 = 970,299 From 5^3 to 99^3 = 95 perfect cubes between 100 and 10,000,000! Am I anywhere close? lol

OpenStudy (sooobored):

uhhh not sure what exactly your problem is asking based off the question and the work provided can you give us any additional information?

OpenStudy (sooobored):

perhaps the units of 100 and 10,000,000?

OpenStudy (mowrey):

The question is what I'm trying to get an answer for. I included what I kind of concluded, but I'm not sure if I was correct.

OpenStudy (sooobored):

between 100 what?

OpenStudy (mowrey):

So I got it or nah? lol

OpenStudy (sooobored):

oh, i see what your question is asking for is 100^3 not a perfect cube?

OpenStudy (sooobored):

so rather than from 5-99 wouldnt it be 5-100?

OpenStudy (sooobored):

and since the question is asking for number of perfect cubes, wouldnt you need to subtract the lowest possible cube from the highest possible cube? or 100-5 = answer?

OpenStudy (sooobored):

all in all, your thought process is correct oh, you did write 95 perfect cubes misleading since you wrote 5^3- 99^3 YES I AGREE WITH 95

OpenStudy (mowrey):

Fantastic! Thanks!